\begin{thebibliography}{10}
\bibitem{asc}
W. Al-Salam,  L. Carlitz,  General Tur\'an expressions for certain
hypergeometric series, Portugal. Math. {\bf 16}, (1957),  119--127.
\bibitem{a1}
R. Askey, Linearization of the product of orthogonal polynomials,
Problems in Analysis, R. Gunning, ed., Princeton University Press,
Princeton, N.J., (1970) 223--228.
\bibitem{bi} J. Bustoz, M. E. H.  Ismail, Tur\'{a}n inequalities for
ultraspherical and continuous $q$--ultraspherical polynomials, SIAM J.
Math. Anal. {\bf 14} (1983), 807--818.
\bibitem{bi1} J. Bustoz, M. E. H.  Ismail,  Tur\'{a}n inequalities for
symmetric orthogonal polynomials, Internat. J. Math. Math. Sci. {\bf 20} (1997), no. 1, 1--7.
\bibitem{ch} T. Chihara, An Introduction to Orthogonal Polynomials, Mathematics and Its Applications, Vol. 13, Gordon 6 Breach, New York, 1978.
\bibitem{g1}
G. Gasper,  On the extension of Tur\'{a}n's inequality to Jacobi
polynomials, Duke. Math. J. {\bf 38} (1971), 415--428.
\bibitem{g2} G. Gasper, An inequality of Tur\'{a}n type for Jacobi
polynomials, Proc. Amer. Math. Soc. {\bf 32}  (1972), 435--439.
\bibitem{ks}
S. Karlin, G. Szeg\"o,  On certain determinants whose elements are
orthogonal polynomials, J. d'Analyse Math. {\bf 8} (1960/61), 1--157.
\bibitem{mane} A. M\'at\'{e}, P. Nevai,  Orthogonal polynomials and absolutely continuous measures, In C. K. Chui et. al., editor, Approximation Theory IV, Vol. 103, Academic Press, New York,   (1983), 611--617.
\bibitem{mn}
B. N. Mukherjee, T. S. Nanjundiah,  On an inequality relating to Laguerre
and Hermite polynomials, Math. Student. {\bf 19} (1951), 47--48.
\bibitem{ne} P. Nevai, Orthogonal Polynomials, Memoirs Amer. Math. Soc. 18
No. 213 (1979).
\bibitem{sk}
H. Skovgaard,  On inequalities of the Tur\'an type, Math.  Scand.  {\bf 2} (1954),
65--73.
\bibitem{szasz}
O. Sz\'asz,  Identities and inequalities concerning orthogonal
polynomials and Bessel functions, J. d'Analyse Math.  {\bf 1} (1951), 116--134.
\bibitem{szego} G. Szeg\"o,   Orthogonal Polynomials, 4th Ed. Amer. Math. Soc. Colloq. Publ. Vol. 23, 1975.
\bibitem{szego1}
G. Szeg\"o On an inequality of P. Tur\'an concerning Legendre
polynomials, Bull. Amer. Math. Soc. {\bf 54} (1948), 401--405.
\bibitem{szego2}
G. Szeg\"o, An inequality for Jacobi polynomials, Studies in Math.
Anal.  and Related Topics, Stanford Univ.  Press, Stanford, Calif.,
(1962), 392--398.
\bibitem{szw0} R. Szwarc,  Orthogonal polynomials and a discrete boundary value problem I, SIAM Journal on Mathematical Analysis {\bf 23} (1992), 959--964.
\bibitem{szw} R.  Szwarc, Positivity of Tur\'{a}n determinants for orthogonal polynomials. In: Harmonic Analysis and Hypergroups, Eds. K.A. Ross et al.,
Birkhauser, Boston-Basel-Berlin, 1998, 165--182
\bibitem{szw1} R. Szwarc,  Orthogonal polynomials and Banach algebras, "Inzell Lectures on Orthogonal Polynomials,", W. zu Castell, F. Filbir, B. Forster (eds.), Advances in the Theory of Special Functions and Orthogonal Polynomials Nova Science Publishers, vol. 2 (2005), 103-139.
\bibitem{tn}
V. R. Thiruvenkatachar, T. S. Nanjundiah,  Inequalities concerning Bessel
functions and orthogonal polynomials, Proc.  Indian Acad.  Sci.  {\bf
33} (1951) 373--384.
\bibitem{tu}
P. Tur\'an, On the zeros of the polynomials of Legendre, \v{C}asopis
\v{P}est. Mat. {\bf 75} (1951) 113--122. 
\end{thebibliography}

\documentclass [12pt]{article}
\usepackage{amssymb,amsmath,amsthm}

\usepackage[cp1250]{inputenc}
\usepackage[active]{srcltx}

\newtheorem{example}{Example}
\newtheorem{lemma}{Lemma}
\newtheorem{rem}{Remark}

\newtheorem{cor}{Corollary}
\newtheorem{theorem}{Theorem}
\def\supp{{\rm supp}\,}

\long\def\comment#1{{}}

\providecommand{\keywords}[1]
{
  \small	
  \textbf{\textit{Keywords---}} #1
}
\begin{document}
\title{Positivity of Tur\'an determinants for orthogonal polynomials II}
\author{Ryszard Szwarc}
\date{}

\maketitle
\begin{abstract}
The polynomials $p_n$ orthogonal on the interval $[-1,1],$ normalized by $p_n(1)=1,$ satisfy Tur\'an's inequality if
$p_n^2(x)-p_{n-1}(x)p_{n+1}(x)\ge 0$ for $n\ge 1$ and for all $x$ in the
interval of orthogonality.  We give a general criterion for orthogonal
polynomials  to satisfy Tur\'an's inequality.  This extends essentially the results of \cite{szw}. In particular the results can be applied to many classes of orthogonal polynomials, by inspecting their recurrence relation.
\end{abstract}

\footnotetext[1]{\noindent 1991 {\it Mathematics Subject Classification}.
Primary  42C05, 47B39}
\keywords{orthogonal polynomials, Tur\'an
determinants, recurrence formula}
\section{Introduction}
Consider a symmetric probability measure $\mu$ such that $\supp\mu= [-1,1].$  By the Gram-Schmidt orthogonalization procedure applied to the system of monomials $x^n,$ $n\ge 0, $ we obtain a sequence of orthogonal polynomials $p_n(x),$ $n\ge 0.$ Every polynomial $p_n$ is of exact degree $n.$ We may assume that its  leading coefficient is positive.
It is well known that the polynomials $p_n$ satisfy the three term recurrence relation of the form
\begin{equation}\label{recrel}
xp_n= \gamma_n p_{n+1}+\alpha_np_{n-1},\quad n\ge 0,
\end{equation} with convention $\alpha_0=p_{-1}=0.$
Due to orthogonality the polynomial $p_n$ has $n$ roots  in the open interval $(-1,1).$ Therefore $p_n(1)>0.$ Let
$$P_n(x)={p_n(x)\over p_n(1)}, \quad n\ge 0.$$
The coefficients $\gamma_n, \ \alpha_{n+1}$ are positive for  $n\ge 0.$ In case the polynomials $p_n$ are orthonormal then the sequences of the coefficients are related by $\gamma_n=\alpha_{n+1}$ and the recurrence relation simplifies to
$$xp_n=\alpha_{n+1}p_{n+1}+\alpha_np_{n-1}, \qquad n\ge 0.
$$
 We refer to \cite{ch,szego} for the basic theory concerning orthogonal polynomials.

We are interested in determining when 
\begin{equation}\label{main}
\Delta_n(x):=P_n(x)^2-P_{n-1}(x)P_{n+1}(x)\ge 0, \qquad n\ge 1.
\end{equation}
The expression $\Delta_n(x)$ is called the Tur\'an's determinant. The problem has been studied for many classes of specific orthogonal polynomials (see \cite{asc, a1, bi, bi1, g1, g2, ks, mn, sk, szasz, szego1, szego2, tn, tu}. We refer to the introduction in \cite{szw} for a short account of known results.

 Tur\'an determinants can be used 
 to determine the orthogonality measure $\mu$ in terms of orthonormal polynomials  $p_n.$
Paul Nevai \cite{ne} observed
if 
  $\alpha_n\stackrel{n}{\to} 1/2$  then the sequence of measures (perhaps signed) 
$$[p_n^2(x)-p_{n-1}(x)p_{n+1}(x)]\,d\mu(x)$$ is weakly convergent to the measure 
$${2\over \pi} \sqrt{1-x^2}\,dx, \quad |x|<1.$$
M\'at\'e and Nevai \cite{mane} showed that if additionally  sequence $\alpha_n$ has bounded variation then the limit of Tur\'an determinants exists. Moreover the orthogonality measure is absolutely continuous on the interval $(-1,1)$ 
its density is given by
 $${2\sqrt{1-x^2}\over \pi f(x)}, \quad |x|<1,$$
 where 
$$f(x):=\lim_n [p_n^2(x)-p_{n-1}(x)p_{n+1}(x)]>0,\quad |x|<1.$$
 
It turns out that the way we normalize the polynomials is essential for
the Tur\'an inequality to hold.
 Indeed, assume  $p_n$ satisfy (\ref{recrel})
and $p_n(1)=1,$ i.e.
\begin{equation}\label{one}
 \alpha_n+\gamma_n=1.
\end{equation} 
Assume
$$p_n^2(x)-p_{n-1}(x)p_{n+1}(x)\ge  0,\quad | x|\le 1,\ n\ge 1.$$
Define new polynomials by $p^{(\sigma)}_n(x)=\sigma_np_n(x),$ where $\sigma_n$ is a sequence of positive coefficients.  Then the condition
$$\{p^{(\sigma)}_n(x)\}^2-p^{(\sigma)}_{n-1}(x)p^{(\sigma)}_{n+1}(x)\ge
0,\quad|x|\le 1,\ n\ge 1$$
is equivalent to (see Proposition \cite{szw})
$$\sigma_n^2 -\sigma_{n-1}\sigma_{n+1}\ge 0,\quad n\ge 1.$$
This means if the Tur\'an determinants are nonnegative, when the polynomials are normalized at $x=1,$ then they stay nonnegative for any other normalization provided that they are nonnegative at $x=1,$ as $\sigma_n=p_n^{(\sigma)}(1).$

By Theorem 1 \cite{szw} if the polynomials are normalized at $x=1,$ i.e. $p_n(1)=1,$ $\alpha_n$ is increasing and $\alpha_n\le {1\over 2},$ the Tur\'an determinants are positive in the interval $(-1,1).$ This result can be applied to many classes of orthogonal polynomials, including for example the ultraspherical polynomials for which positivity has been obtained in 
\cite{sk, szasz} 

The result mentioned above can be applied provided that we are given the coefficients $\alpha_n$ explicitly. 
For many classes of orthogonal polynomials in the interval $[-1,1]$ we are given recurrence relations, but the values $p_n(1)$ cannot be evaluated in the explicit form. Therefore we are unable to provide a recurrence relation  for the polynomials $P_n(x)=p_n(x)/p_n(1),$ in the form for which we can inspect easily the assumptions of Theorem 1 \cite{szw}. This occurs when
we study the associated polynomials. Indeed assume $p_n$ satisfy (\ref{recrel}) and (\ref{one}). For a fixed natural number the associated polynomials $p_n^{(k)}$ of order $k$ are defined
by
\begin{equation}
xp_n^{(k)}=
\begin{cases}
\gamma_kp_1^{(k)} & n=0,\\
\gamma_{n+k}p_{n+1}^{(k)}+\alpha_{n+k}p_{n-1}^{(k)} & n\ge 1.
\end{cases}
\end{equation}
These polynomials do not satisfy $p_n^{(k)}(1)=1$ as
$$p_1^{(k)}(1)=\gamma_k^{-1}=(1-\alpha_k)^{-1}>1.$$

The obstacle described above has been partially overcome in Corollary 1 of \cite{szw}, but it required additional assumptions, in particular $\gamma_0\ge 1.$ 
Unfortunately many examples including the associated polynomials violate that condition. 
The aim of this note is to provide a counterpart to Corollory 1 \cite{szw}
by allowing $\gamma_0<1.$ This is done in Theorem 1. As the assumptions in this theorem are complicated Corollary 1 provides a wide class of relatively simple recurrence relations for which Theorem 1 applies. General examples are provided at the end of the paper.
\section{Results}
\begin{theorem}
Assume the polynomials $p_n$ satisfy
\begin{equation}\label{1}
xp_n=\gamma_np_{n+1}+\alpha_np_{n-1},\quad n\ge 0,
\end{equation} where $\alpha_0=p_{-1}=0,\ p_0=1.$ 
Assume 
\begin{enumerate}
\item[(a)] the sequence $\alpha_{n}$ is strictly increasing and $\alpha_{n}\le {1/2},$
\item[(b)] the sequence $\gamma_n$ is positive and strictly decreasing ,
\item[(c)] $\alpha_n+\gamma_n\le 1.$
\end{enumerate}
Assume also that  there holds
\begin{eqnarray}
{\alpha_n-\alpha_{n-1}\over \alpha_{n}\gamma_{n-1}-\alpha_{n-1}\gamma_{n}} &\le & 
{\alpha_{n+1}\gamma_n-\alpha_n\gamma_{n+1}\over \gamma_n-\gamma_{n+1}}, \quad n\ge 1,\label{2}\\
\gamma_0-\gamma_1&\le & \alpha_1\gamma_0^2.\label{3}
\end{eqnarray}
 Then for $${P}_n(x)={p_n(x)\over p_n(1)}$$ we have
 $${P}_n(x)^2-{P}_{n-1}(x){P}_{n+1}(x)\ge 0,\qquad -1\le x\le 1.$$
\end{theorem}
\begin{proof}
Let $$g_n={p_{n+1}(1)\over p_n(1)}.$$
By (\ref{1}) we get
\begin{equation}\label{5}
g_n={1\over \gamma_n}\left (1-{\alpha_n\over g_{n-1}}\right ), \quad n\ge 1.
\end{equation}
\begin{lemma} Under assumptions of Theorem 1 there holds
\begin{equation}\label{6}
1\le g_n\le {\alpha_{n+1}\gamma_n-\alpha_n\gamma_{n+1}\over \gamma_n-\gamma_{n+1}}, \quad n\ge 0.
\end{equation}
\end{lemma}
\begin{proof}
(\ref{1}) gives $g_0=1/\gamma_0\ge 1.$ Assume $g_{n-1}\ge 1$ for $n\ge 1.$ By (\ref{5}) and (c) we get
$$g_n\ge {1\over \gamma_n}\left (1-\alpha_n\right )\ge 1.$$ This shows the left hand side inequality. 

By (\ref{3}) we get
$$g_0={1\over \gamma_0}\le {\alpha_1\gamma_0\over \gamma_0-\gamma_1},$$
which shows the right hand side inequality in (\ref{6}) for $n=0.$ Assume (\ref{6}) holds for some $n\ge 0.$ Then, in view of (\ref{5}) and (\ref{2}), we get
\begin{multline*}
g_{n+1}={1\over \gamma_{n+1}}\left (1-{\alpha_{n+1}\over g_n}\right )\le 
{1\over \gamma_{n+1}}\left ( 1-{\alpha_{n+1}(\gamma_n-\gamma_{n+1})\over \alpha_{n+1}\gamma_n-\alpha_n\gamma_{n+1}}\right )\\
= {\alpha_{n+1}-\alpha_n\over \alpha_{n+1}\gamma_n-\alpha_n\gamma_{n+1}}\le 
{\alpha_{n+2}\gamma_{n+1}-\alpha_{n+1}\gamma_{n+2}\over \gamma_{n+1}-\gamma_{n+2}}.
\end{multline*}
\end{proof}
\begin{lemma}
Under the assumptions of Theorem 1 the sequence $g_n=p_{n+1}(1)/p_n(1)$ is nonincreasing.
\end{lemma}
\begin{proof}
Let
$$f_k(x)={1\over \gamma_k}\left (1-{\alpha_k\over x}\right ),\quad x\ge 1.$$ The functions $f_k$ are nondecreasing. Moreover by a straightforward computation we get
\begin{equation}\label{f}
f_{k+1}(x)\le f_k(x),\qquad 1\le x\le {\alpha_{k+1}\gamma_k-\alpha_k\gamma_{k+1}\over \gamma_k-\gamma_{k+1}}.
\end{equation}
We have $$g_0={1\over \gamma_0},\quad g_1={1\over \gamma_1}(1-\alpha_1\gamma_0).$$ By (\ref{3}) we get $g_0\ge g_1.$ Assume $g_{n-1}\ge g_n.$
Then in view of (\ref{5}) and Lemma 1 we obtain
$$g_{n+1}=f_{n+1}(g_n)\le f_n(g_n)\le f_n(g_{n-1})=g_n.$$
\end{proof}
The polynomials ${P}_n$ satisfy
$$x{P}_n=\widetilde{\gamma}_n{P}_{n+1}+
\widetilde{\alpha}_n{P}_{n-1},\quad n\ge 0,$$
where
$$\widetilde{\alpha}_n=\alpha_n {p_{n-1}(1)\over p_n(1)},\qquad \widetilde{\gamma}_n=\gamma_n {p_{n+1}(1)\over p_n(1)}.$$
Since ${P}_n(1)=1$ we get
$$ \widetilde{\alpha}_n+\widetilde{\gamma}_n=1.$$ Moreover by Lemma 1,  Lemma 2 and (a)  the sequence 
$\widetilde{\alpha}_n$ is nondecreasing and 
$\widetilde{\alpha}_n\le 1/2.$ Thus the conclusion follows from Theorem 1(i) of \cite{szw}.
\end{proof}
\begin{rem} {\rm
As a side effect of Theorem 1 we get that the polynomials $p_n$ admit nonnegative linearization as the polynomials ${P}_n$ satisfy the assumptions of Theorem 1 in \cite{szw0}. We refer to \cite{szw1} where this  problem is discussed in detail.}
\end{rem}

The assumption (6) in Theorem 1 can be troublesome for verification in examples. However there is a wide class of examples for which (6) simplifies substantially. 
\begin{cor}
Let the polynomials $p_n$ satify (\ref{1}) with
$$\alpha_n={1\over 2}-\alpha \delta_n, \qquad \gamma_n={1\over 2} +\gamma \delta_n,\quad n\ge 0.$$ where
$\alpha\ge \gamma> 0$ and $\delta_n\searrow 0.$ 
Then the conclusion of Theorem 1 holds. 
\end{cor}
\begin{proof}
We have
$$ \alpha_{n+1}\gamma_n-\alpha_n\gamma_{n+1}={1\over 2}(\alpha+\gamma)(\delta_n-\delta_{n+1}), \quad n\ge 0.$$ Thus (\ref{2}) takes the form
$${2\alpha\over \alpha+\gamma}\le {\alpha+\gamma\over 2\gamma},$$ which is true for any  numbers $\alpha, \gamma >0.$

Next, since  $$0=\alpha_0={1\over 2}-\alpha\delta_0$$ we get $\alpha\delta_0= 1/2.$ Thus 
$$
\alpha_1\gamma_0^2=\left ({1\over 2}-{1\over 2} {\delta_1\over \delta_0}\right )\left ({1\over 2} +\gamma\delta_0\right )^2\ge {\delta_0-\delta_1\over 2\delta_0} \,2\gamma\delta_0=\gamma_0-\gamma_1.
$$
Therefore all the assumptions of Theorem 1 are satisfied.
\end{proof}
\begin{example}{\rm 
Consider the symmetric Pollaczek polynomials
$P_n^\lambda(x;a).$ They are orthogonal in the interval $[-1,1]$
and satisfy the recurrence relation
$$
xP_n^\lambda(x;a)= {n+1\over 2(n+\lambda+a)}
P_{n+1}^\lambda(x;a) + {n+2\lambda -1\over 2(n+\lambda
+a)}P_{n-1}^\lambda(x;a),$$
where the parameters satisfy $a>0,\ \lambda >0.$
Set
$$p_n(x)= {n!\over (2\lambda)_n}P_n^\lambda(x;a),$$
where $(\mu)_n=\mu(\mu+1)\ldots (\mu +n-1).$
Then the polynomials $p_n$ satisfy the recurrence relation
$$xp_n=
{n+2\lambda\over 2(n+\lambda+a)}p_{n+1}+
{n\over 2(n+\lambda+a)}p_{n-1}.$$
Observe that the  assumptions of
Corollary 1(i) of \cite{szw} are satisfied for $a\ge \lambda.$ 

\begin{rem}\rm  There is a misprint  in the formulation of Corollary 1 in \cite{szw}. The assumptions there required that
$$\lim_n\alpha_n ={1\over 2}\widetilde{a},\qquad \lim_n\gamma_n={1\over 2}\widetilde{a}^{-1}$$ with $0<\widetilde{a}<1.$ But the conclusion holds also for $\widetilde{a}=1$ with the same proof as in \cite{szw}. For symmetric Pollaczek polynomials we actually have $\widetilde{a}=1.$
\end{rem}
However for $\lambda>a$ the assumptions  of Corollary 1(ii) \cite{szw} are not satisfied  as was wrongly stated in \cite{szw}, because $\gamma_0<1.$  Instead we can apply  Corollary 1, on the previous page, with
$$\alpha=\lambda+a,\quad \gamma =\lambda-a,\quad \delta_n={1\over 2(n+\lambda+a)}.$$}
\end{example}
\begin{rem}{\rm Corollary 1 requires $\alpha\delta_0={1\over 2},$ i.e. the quantity $\delta_0$ is determined by $\alpha,$ which limits the range of examples. We will  get rid of that  assumption in the next corollary, allowing some flexibility for the quantity $\delta_0$. }
\end{rem}
\begin{cor}
Let the polynomials $p_n$ satify (\ref{1}) with
$$\alpha_0=0, \qquad \gamma_0={1\over 2}+\gamma\delta_0,$$
$$\alpha_n={1\over 2}-\alpha \delta_n, \qquad \gamma_n={1\over 2} +\gamma \delta_n,\quad n\ge 1,$$ where
$\alpha\ge \gamma> 0$ and $\delta_n\searrow 0.$ Assume also that
\begin{equation}\label{strange}{3\gamma-\alpha\over 2\gamma(\alpha+\gamma)}\le\delta_0\le  {1\over 2\alpha}.
\end{equation}
Then the conclusion of Corollary 1 holds. 
\end{cor}

\begin{rem}{\rm 
The condition $\delta_0\le 1/(2\alpha)$ is not artificial. Instead of setting $\alpha_0=0$ we could define
$$\alpha_0={1\over 2}-\alpha\delta_0.$$ The aformentioned assumption amounts to the condition $\alpha_0\ge 0.$

Observe also that the possible range for the quantity $\delta_0$ described in (\ref{strange}) is nonempty as
we always have
$${3\gamma-\alpha\over 2\gamma(\alpha+\gamma)}\le  {1\over 2\alpha}.$$}
\end{rem}
\begin{proof}
We are forced to modify the proof of the preceding corollary at  places where $\delta_0$ shows up, as $\alpha_0=0$ is no longer equal ${1\over 2}-\alpha\delta_0.$ Thus we have to make calculations concerning   (\ref{2}), for $n=1,$  and (\ref{3}), by hand.  Since $\alpha\delta_0\le {1\over 2}$ we get
$$\alpha_1\gamma_0^2\ge \left ({1\over 2}-{1\over 2} {\delta_1\over \delta_0}\right )\left ({1\over 2} +\gamma\delta_0\right )^2\ge {\delta_0-\delta_1\over 2\delta_0} \,2\gamma\delta_0=\gamma_0-\gamma_1.
$$ This gives (\ref{3}).
Next  we verify (\ref{2}) for $n=1,$ as the value $\delta_0$ is involved there on the left hand side. The inequality (\ref{2}) in this case reduces to
$${1\over \gamma_0}={2\over 1+2\gamma\delta_0}\le {\alpha+\gamma\over 2\gamma}.$$
This inequality is equivalent to the left hand side of (\ref{strange}).
\end{proof}
\begin{rem}\label{long}
\rm Corollary 1 requires that the sequence
\begin{equation}\label{7}
{\gamma_n-{1\over 2}\over {1\over 2}-\alpha_n}, \quad n\ge 1
\end{equation} is constant.
It is possible to  extend Corollary 1 to the case when the sequence in (\ref{7}) is  nondecreasing. Indeed
\begin{multline}\label{last}
\alpha_{n+1}\gamma_n-\alpha_n\gamma_{n+1} =\left [\left (\gamma_{n+1}-\textstyle{1\over 2}\right )\left (\textstyle{1\over 2}-\alpha_n\right )-\left (\gamma_{n}-\textstyle{1\over 2}\right )\left (\textstyle{1\over 2}-\alpha_{n+1}\right )\right ]
\\ +\textstyle{1\over 2}(\alpha_{n+1}-\alpha_n +\gamma_n-\gamma_{n+1}) \ge \textstyle {1\over 2}(\alpha_{n+1}-\alpha_n +\gamma_n-\gamma_{n+1}).
\end{multline}
Denote
\begin{equation}\label{uv}
u_n=\alpha_{n+1}-\alpha_n,\quad v_n=\gamma_n-\gamma_{n+1}.
\end{equation} By (\ref{last}) the assumption (\ref{2}) will be satisfied if
\begin{equation}\label{8}
(u_{n-1}+v_{n-1})(u_n+v_n)\ge 4u_{n-1}v_n.
\end{equation}
Let 
\begin{equation}\label{lambda}
v_k=\lambda_ku_k, \quad 0<\lambda_k\le 1.
\end{equation} Then (\ref{8}) takes the form
$$
(1+\lambda_{n-1})(1+\lambda_n)\ge 4\lambda_n
$$
i.e.
\begin{equation}\label{9}
\lambda_n\le {1+\lambda_{n-1}\over 3-\lambda_{n-1}}.
\end{equation}
Let
$$f(x)={1+x\over 3-x},\quad 0\le x\le 1.$$ The condition (\ref{9}) amounts to 
\begin{equation}\label{f}
\lambda_n\le f(\lambda_{n-1}).
\end{equation}
Thus (\ref{f}) implies (\ref{2}), provided that
the sequence in (\ref{7}) is nondecreasing.
As $f(x)\ge {1\over 3},$ the inequality
 (\ref{f}), and consequently (\ref{2}),  is satisfied whenever $\lambda_n\le 1/3.$
 Observe that for $y\ge 1$ we have
\begin{equation}\label{10}
 f\left ({y-1\over y+1}\right )={y\over y+2}.
\end{equation}
\end{rem}
Remark \ref{long} gives rise to new examples.
\begin{example}\rm  For $\varepsilon_n\searrow 0,$ $\delta_n\searrow \delta\ge 0,$ let
$$\alpha_n={1\over 2}-3\varepsilon_n(1+\delta_n),\qquad \gamma_n={1\over 2}+{ \varepsilon_n}, \qquad n\ge 0.$$
Then
$${\gamma_n-{1\over 2}\over {1\over 2}-\alpha_n}={1\over 3(1+\delta_n)}\nearrow {1\over 3(1+\delta)}$$ and (see (\ref{uv}) and (\ref{lambda}))
$$\lambda_n={\varepsilon_n-\varepsilon_{n+1}\over {3(\varepsilon_n-\varepsilon_{n+1})+3(\varepsilon_n\delta_n-\varepsilon_{n+1}\delta_{n+1})}}\le {1\over 3}.$$
Next
$$1+\delta_1\le 1+\delta_0={1\over 6\varepsilon_0}$$ (the last equality follows from $\alpha_0=0$).
Then
\begin{multline*}
\alpha_1\gamma_0^2=\left [{1\over 2}-3\varepsilon_1(1+\delta_1)\right ]\,\left ({1\over 2}+\varepsilon_0\right )^2 \\ \ge {1\over 2}\left ( 1-{\varepsilon_1\over \varepsilon_0}\right )\left ({1\over 2}+\varepsilon_0\right )^2
\ge {1\over 2}\left (1-{\varepsilon_1\over \varepsilon_0} \right )\, 2\varepsilon_0=\gamma_0-\gamma_1.
\end{multline*}
This gives (\ref{3}).
\end{example}

\begin{example}\label{exm}
\rm  For $a>0$  let
$$\alpha_n={1\over 2}-{a\over 2(n+a)},\qquad
\gamma_n={1\over 2} +{a\over 2(n+a+1)}.$$ Then the sequence in (\ref{7}) is increasing. Furthermore (cf. (\ref{uv}) and (\ref{lambda}))
$$ u_n={a\over 2(n+a)(n+a+1)},\quad v_n={a\over 2(n+a+1)(n+a+2)},\quad \lambda_n= {n+a\over n+a+2}.$$
By (\ref{10}) we have
$f(\lambda_{n-1}) =\lambda_n.$ Thus   (\ref{2}) is satisfied.   Next
\begin{align*}
&\gamma_0-\gamma_1=v_0={a\over 2(a+1)(a+2)}\le {a\over 2(a+1)^2}, \\ &\alpha_1\gamma_0^2={(2a+1)^2\over 8(a+1)^3}.
\end{align*}
As
$$
(2a+1)^2\ge 4(a+1)a,
$$ we get
$$\alpha_1\gamma_0^2\ge {a\over 2(a+1)^2}\ge \gamma_0-\gamma_1,$$
so
the condition (\ref{3}) is also satisfied.
\end{example}

\begin{rem}\label{rm}\rm
\ Let $$\lambda_n={y_n-1\over y_n+1}.$$  Then
\begin{equation}\label{y}
y_n={1+\lambda_n\over 1-\lambda_n}.
\end{equation}
 Moreover condition (\ref{9}) is equivalent to
\begin{equation}\label{in}
y_n\le y_{n-1}+1.
\end{equation}
\end{rem}

Using Remark \ref{rm} we can still generalize Example \ref{exm}.
\begin{example}\rm  For $a >0, \ b\ge 0$  let
$$\alpha_n={1\over 2}-{a\over 2(n+a)},\qquad
\gamma_n={1\over 2} +{a\over 2(n+a+b+1)}.$$  The sequence in (\ref{7}) is increasing. Next
\begin{align*}
&u_n={a\over 2(n+a)(n+a+1)},\\
& v_n={a\over 2(n+a+b+1)(n+a+b+2)},\\ &\lambda_n= {(n+a)(n+a+1)\over (n+a+b+1)(n+a+b+2)}.
\end{align*} 
By (\ref{y}) we get
$$y_n={n\over b+1}+{2a+b+2\over 2(b+1)}+{b^2+2b\over 2(b+1)(2n+2a+b+2)}.$$
Since $b\ge 0,$ the inequality (\ref{in}) holds. Next
\begin{align*}
&\gamma_0-\gamma_1=v_0= {a\over 2(a+b+1)(a+b+2)}\le {a\over 2(a+b+1)^2},\\
&\alpha_1\gamma_0^2={(2a+b+1)^2\over 8(a+1)(a+b+1)^2}\ge {(2a+1)^2\over 8(a+1)(a+b+1)^2} \ge {a\over 2(a+b+1)^2} .
\end{align*}
Thus
 (\ref{3}) is fulfilled.
\end{example}

{\bf Acknowledgment} I am grateful to the referees. Their remarks improved the exposition substantially.

\obeylines{
Institute of Mathematics
Wroc{\l}aw University
pl. Grunwaldzki 2/4
50--384 Wroc{\l}aw, Poland
{\tt ryszard.szwarc@uwr.edu.pl}}
\end{document}